# HalluciBot: Is There No Such Thing as a Bad Question?

### v1

@@ Abstract
Hallucination continues to be one of the most critical challenges in the institutional adoption journey of Large Language Models (LLMs). In this context, an overwhelming number of studies have focused on analyzing the post-generation phase - refining outputs via feedback, analyzing logit output values, or deriving clues via the outputs’ artifacts. We propose HalluciBot, a model that predicts the probability of hallucination before generation, for any query imposed to an LLM. In essence, HalluciBot does not invoke any generation during inference. To derive empirical evidence for HalluciBot, we employ a Multi-Agent Monte Carlo Simulation using a Query Perturbator to craft $n$ variations per query at train time. The construction of our Query Perturbator is motivated by our introduction of a new definition of hallucination - truthful hallucination. Our training methodology generated 2,219,022 estimates for a training corpus of 369,837 queries, spanning 13 diverse datasets and 3 question-answering scenarios. HalluciBot predicts both binary and multi-class probabilities of hallucination, enabling a means to judge the query’s quality with regards to its propensity to hallucinate. Therefore, HalluciBot paves the way to revise or cancel a query before generation and the ensuing computational waste. Moreover, it provides a lucid means to measure user accountability for hallucinatory queries.
HalluciBot: Is There No Such Thing as a Bad Question?
William Watsonthanks: Equal Contributionand Nicole Cho [^1] Equal Contribution J.P. Morgan AI Research New York, NY, USA william.watson@jpmchase.com nicole.cho@jpmorgan.com
@@ 1 Introduction
Figure 1: Comparison between traditional methods of confidence estimation and HalluciBot. In the former, user inputs a query to generate an output that is accurate or hallucinatory. If hallucinatory, user can end the session or revise the query for iterative rounds of generation. In contrast, HalluciBot predicts the probability of hallucination for a given query - before generation. Therefore, users can instantly gain insight into hallucination probability and revise a query.
Despite the promising potential for a myriad of practical use cases, Large Language Models (LLMs) offer limited insights into their chain of thought (Liang et al., 2022; Wei et al., 2023; Kojima et al., 2023; Li et al., 2023) and have the propensity to hallucinate in various circumstances (Jiang et al., 2021). Common factors that drive hallucinations encompass high model complexity, flawed data sources, or inherent sampling randomness. Specifically, the intrinsic trade-off between greedy deterministic decoding and the creativity spawned through nucleus sampling induces a heightened propensity to generate hallucinations (Huang et al., 2023). This challenge is compounded by limitations such as the frequent inaccessibility into the LLMs’ training datasets (Liang et al., 2022). HuggingFace’s release of its "Hallucinations Leaderboard" on January 29th, 2024 (Minervini et al., 2024; Gao et al., 2023) highlights the importance of resolving hallucination-related issues via the concerted effort of evaluating different LLMs. In this context, the majority of current studies have focused on the post-generation phase of output analysis as expanded in Peng et al. (2023) such as - (1) self-refinement via feedback loops on the model’s output (Madaan et al., 2023), (2) analysis of logit output values to detect hallucination (Varshney et al., 2023), or (3) for a minority of studies focused on the pre-generation phase, the ingestion of recent knowledge to improve performance (Tonmoy et al., 2024). We propose a novel model, HalluciBot, that predicts the probability of hallucination, before any generation, for any type of query. In essence, for any query posed to an LLM, HalluciBot estimates:
Figure 2: System Overview. A single query $q_{0}$, supplied by the user, is perturbed in $n$ different ways. Next, the original and perturbed queries $q_{i}$ are independently answered by the Generator agents. This Monte Carlo Simulation provides an estimate into the rate of hallucination $p_{h}(q_{0})$ for an original query $q_{0}$. Undergoing this Multi-Agent Monte Carlo Simulation, covering 13 datasets, HalluciBot is trained to predict the probability that any query $q_{0}$ will yield an hallucination, and predict the expected value of hallucinations sampled - before generation.
@@
    a binary classification of the query’s propensity to hallucinate ("Yes" or "No"), as well as,
@@
    a multi-class hallucination rate estimating the expected value of hallucination via sampling.
In order to understand HalluciBot, it is imperative to discuss the following key fundamental characteristics regarding hallucinations and LLMs. Firstly, hallucination itself is not a derivative of any mathematical model and intrinsically, can only be determined by juxtaposing the ground truth with the LLM’s generated outputs. Secondly, LLMs frequently advance output quality by different sampling methods such as nucleus sampling (Holtzman et al., 2020; Fan et al., 2018; Holtzman et al., 2018; Radford et al., 2018). In light of these features, HalluciBot is trained via the method of leveraging a Multi-Agent Monte Carlo Simulation for the construction of an encoder for hallucination quantification. In essence, HalluciBot, leveraging gpt-3.5-turbo, is trained via:
[…]
    perturbing 369,837 queries $n$ times into unique variations that retain the original semantic meaning yet significantly diverge lexically; then,
[…]
    employing $n+1$ independent agents to sample an output from each query (including the original) at a temperature of $1.0$ for diversity;
[…]
    conducting a Monte Carlo Simulation on 2,219,022 sampled outputs;
[…]
    deriving an empirical estimate into the expected rate of hallucination $p_{h}(q_{0})$ for the original query as the ground truth for our encoder.
As a result, our study has culminated in the following pillars of contribution.
[…]
    HalluciBot, the first encoder-based model to derive, before generation, an anticipated rate of hallucination for any type of query. HalluciBot has scored a validation accuracy of 73% on binary labels and 73% for top 3 multi-class labels.
[…]
    Moreover, our approach to construct HalluciBot absorbs the computational complexity of Monte Carlo sampling and training prior to the user session. This differs from the prevalent hallucination detection methods that involve sampling during or after the user’s current session. Thus, institutions that employ HalluciBot can systematically save on the considerable amount of computational waste engendered by "highly probable" hallucinatory queries.
[…]
    HalluciBot also generalizes to systems with Retrieval Augmented Generation (RAG) context or few-shot question answering systems with an LLM generator. Also it can generalize to systems only accessible via API calls (OpenAI, 2022; Google, 2023; Microsoft, 2023). Thus, in diverse implementation environments, HalluciBot can impart accountability to the user with regards to the query’s quality.
[…]
    We also present a myriad of behavioral and statistical discoveries from our experiments, such as an LLM’s probability to either be exceptionally correct or completely incorrect in certain scenarios. Our Appendix elaborates in detail.
@@ 2 Related Work
With regards to hallucination mitigation studies, an overwhelming majority focuses on the post-generation stage of analyzing outputs. A minority concentrates on the pre-generation phase and even amongst those, the focus lies in incorporating recent knowledge into LLMs. In detail, many expand on the universally utilized method of context-based retrieval systems (Reimers and Gurevych, 2019; Johnson et al., 2019; Nogueira and Cho, 2020; Karpukhin et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021). Other methods include relying on the model’s general knowledge (Khashabi et al., 2020) or conditioning the QA model on context generated by the LLM itself (Yu et al., 2023). Certain work has focused on mitigating hallucinations by augmenting the way LLMs generate their answers. One of the more popular techniques is to have the model enumerate its chain-of-thought (Wei et al., 2023) and think step by step (Nye et al., 2021), while building context. Another method to augment generation with context is by semantic retrieval (Lewis et al., 2021; Liu et al., 2021), handling hallucinations as they arise (Varshney et al., 2023), or using LLMs to generate context rather than retrieve (Yu et al., 2023). PromptChainer (Wu et al., 2022) profiled techniques to craft LLM chains, in which the output of one LLM’s generation process, when fed into the next LLM, can allow for more complex tasks. Language Model Cascades (Dohan et al., 2022) demonstrated that LLMs can yield probabilistic programs to tackle multi-step reasoning problems. Self-consistency (Wang et al., 2023) leveraged a new decoding strategy to sample multiple generative pathways - then select the most consistent answer. Also, Kumar et al. (2022) explored gradient-based sampling procedures that satisfy user-defined constraints. Most recent work has focused on sampling-based calibration within a single model (Cole et al., 2023) or self-verification (Kadavath et al., 2022) - the latter focuses on generating a set of outputs and feeding those back into the LLM. Furthermore, Snyder et al. (2023) explores how artifacts can differentiate hallucinated outputs. One common feature amongst these approaches is that the focus is on the output rather than the query. Also, while Zheng and Saparov (2023) study lexical perturbations, no study on hallucinations has employed a Multi-Agent approach coupled with query perturbations - which are hallmark features of HalluciBot.
3 Methodology Overview
3.1 What is Hallucination?
In general terms, hallucination refers to a false perception of patterns or objects resulting from one’s senses. With regards to LLMs, a myriad of studies have categorized hallucinations into bifurcated structures such as (1) intrinsic hallucination, which refers to the LLM’s outputs directly contradicting with the source content for extractive queries, or (2) extrinsic hallucination, which refers to the LLM’s outputs being unverifiable by the source content (i.e. irrelevant outputs) Ji et al. (2023). From a different angle, (1) factuality hallucinations refer to outputs which directly contradict or fabricate the ground truth while (2) faithfulness hallucinations define outputs that misunderstand the context or intent of the query (Huang et al., 2023).
In this study, we will focus on the aforementioned types of hallucinations and introduce truthful hallucination as a novel concept.
[…]
    The motivation for truthful hallucination stems from the analysis that neural networks display an intrinsic propensity to memorize training data (Carlini et al., 2021).
[…]
    Truthful hallucination is defined as an LLM’s inability to answer semantically similar but lexically different perturbations of a query.
[…]
    Understanding truthful hallucination is essential to account for training bias. Namely, if an LLM has memorized a query and therefore can answer it correctly, but demonstrates the inability to answer similar queries that are worded differently, it will be held accountable.
[…]
    Moreover, truthful hallucination is reflective of user behavior - unless there is a drop-down of query templates, there will almost always be diverse representations of the same query.
3.2 What is the Motivation for HalluciBot?
HalluciBot focuses on distilling LLM behavior into a speedy, encoder-based system that can predict hallucination before generation. Foremost, this is in contrast to prior work that uses multiple generations during a user’s session to provide self-consistency (Manakul et al., 2023). Next, our proposal differs from entropy based, log-prob based, or model based estimation techniques (Huang et al., 2023) that rely on the LLM’s uncertainty to predict hallucinations - these methods focus on the model’s bias while we focus on empirical estimates. Moreover, our approach consists of a Multi-Agent Simulation which stands in stark contrast to the majority of current experiments that have focused on leveraging a single LLM agent to generate outputs from a single query (Cole et al., 2023; Kadavath et al., 2022; Snyder et al., 2023). The training procedure for HalluciBot consists of:
[…]
    perturbing 369,837 queries $n$ times (§3.3)
[…]
    employing $n+1$ independent LLM agents (§3.4)
[…]
    sampling an output from each query (§3.4)
[…]
    conducting a Monte Carlo Simulation on the 2,219,022 sampled outputs (§3.5)
[…]
    training an encoder-based classifier (§3.6)
3.3 What is a Query Perturbator?
We are motivated by how perturbations:
[…]
    help comprehend truthful hallucination. We induce diversity to disentangle the generation process from any potential training bias.
[…]
    emulate real-life scenarios, in which users unintentionally create variations of the same query.
Thus, the Query Perturbator is a gpt-3.5-turbo LLM agent $\mathcal{T}$ that generates $n=5$ perturbations to the original query $q_{0}$ while retaining the same semantic meaning. Note that in practice, there are $n+1=6$ transformations because for a given query $q_{0}$, we define $T_{0}(q_{0})=I(q_{0})=q_{0}$ as an identity transform. Therefore, the original query will always be included in the set of transformed queries for the next step. In effect, the generation process can be summarized as returning a set of $\mathcal{Q}=\left\{q_{0},q_{1},\ldots,q_{n}\right\}$ query perturbations of size $n+1$. The Query Perturbator’s singular purpose is to: Rewrite the query in {$n$} radically different ways. One prompt call is sufficient to discourage duplicates. Temperature is set to $1.0$ to prioritize creativity and lexical diversity (Table 6).
$\mathcal{Q}=\mathcal{T}(q_{0})=\left[\begin{array}[]{c}I(q_{0})\\ T_{1}(q_{0})\\ \vdots\\ T_{n}(q_{0})\end{array}\right]=\left[\begin{array}[]{c}q_{0}\\ q_{1}\\ \vdots\\ q_{n}\end{array}\right]$
3.4 What is an Output Generator?
Next, the Output Generator is deployed for each perturbation due to the following reasons:
[…]
    Given the complexity of LLMs with nucleus sampling decoding, analyzing a single output from a single query may not be comprehensive enough to determine hallucination.
[…]
    Furthermore, given the risk of over-training LLMs, their opaque training data, and propensity to memorize, generating multiple outputs from the same query does not eliminate training bias. This has been a prevalent mode of studying hallucination in prior studies (§3.2)
A new generation process should be launched each time to answer the next perturbed query to ensure a truly independent output generation process. Thus, for the perturbed set $\mathcal{Q}$ for a sample $q_{0}$, the Output Generator consists of $|\mathcal{Q}|=n+1$ independent gpt-3.5-turbo LLM agents to generate outputs $a_{i}\in\mathcal{A}$ for each variation $q_{i}\in\mathcal{Q}$. The LLM agent will receive (1) for Extractive queries, a prompt with the query $q_{i}$, alongside context $c_{i}$, (2) for Multiple-Choice queries, candidate choices $k_{i}\in\mathcal{K}$, and (3) for Abstractive queries, no additional context. See Table 8 for each experiment’s prompt procedure. Temperature for all experiments is set to $1.0$ to stress-test and encourage diversity.
3.5 What is the purpose of a Monte Carlo Simulation in HalluciBot?
As evidenced by multiple studies and Section 3.1, hallucination is the outcome of multiple confounding variables - thus, it is highly unlikely that a tractable closed-form solution will be able to model hallucinations. Thus, we employ a Monte Carlo Simulation as a means to derive empirical estimations of hallucination rates in LLMs, since this method is frequently leveraged to map probability in the presence of random variable inference (Swaminathan, 2021). Thus, we estimate the probability density that a query induces hallucination.
3.6 How Do We Train a Classifier?
Once the Monte Carlo Simulation is complete for our training corpus composed of 369,837 queries spanning 13 different datasets (Appendix C, Tables 1 & 11), we start training our classifier. These queries encompass Extractive, Multiple Choice, and Abstractive scenarios. Each scenario, with or without additional context, affects the the hallucination rate of gpt-3.5-turbo. The estimates provided through our Monte Carlo Simulation are proportional to an approximation of hallucination rates.
[…]
    With a synthetic labeled set of queries $q_{0}$ and their rate of hallucinations $p_{h}(q_{0})$, we train an encoder-style classifier based on BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019) to estimate the hallucination probability density from our Monte Carlo Simulation.
[…]
    We train two versions: a binary model to estimate the propensity a query can hallucinate, and a multi-class model to predict the expected value of hallucinations if sampled $n+1$ times.
Our experiments constrain the number of perturbations to $n=5$, and when including the original query and output, we can model the hallucination rate for $n+1=6$ modes - this translates to increments of $16.6\%$ in hallucination rates.
Figure 3: Distribution of the expected value (number of hallucinations per example), rendered in percentage. For Extractive, additional context helps mitigate the rate of hallucination. For Multiple Choice, distractors can cause confusion amongst agents uniformly. However, for Abstractive, no additional information can cause massive disparities in correctness - most of our simulations resulted in no hallucinations or all hallucinations.
@@ 3.6 How Do We Train a Classifier?
Extractive (1) SQuADv2
Multiple TruthfulQA, SciQ, MMLU, PIQA,
Choice (9) BoolQ, OpenBookQA, MathQA
ARC - Easy, ARC - Challenge
Abstractive (6) SQuADv2, TruthfulQA, SciQ,
@@
3.7 How Do We Encode a Query’s Scenario?
We conduct an ablation study to explore if incorporating the query’s scenario mitigates hallucinations. To create the prompt, we prepend the original query $q_{0}$ with either [EXTRACTIVE], [MULTIPLE CHOICE], or [ABSTRACTIVE], using the format <<{tag} $\mathtt{q_{0}}$}>>. Our hypothesis is based on recent research that highlights the use of RAG Yan et al. (2024); Lewis et al. (2021); Guu et al. (2020) to alleviate hallucinations. The additional context provides valuable signals related to the hallucination rate of the original query. Furthermore, we apply this technique to distinguish our experimental results from reused datasets in different scenarios, such as SciQ and SQuADv2. The queries are encoded exactly as they appear in the original datasets.
3.8 How Do We Convert Monte Carlo Estimates to Labels?
Empirical Estimate
The probability of truthful hallucinations for a query $q_{0}$, denoted as $p_{h}(q_{0})$, can be empirically estimated based on the output $a_{i}\in\mathcal{A}$ of our Multi-Agent Monte Carlo Simulation. We define the indicator function $\mathbb{I}$ to measure the incorrectness of an output $a_{i}$ with respect to the ground truth $y$ for query $q_{0}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\big{[}a_{i}\not=y\big{]}$
Binary Labels
To assess the propensity to hallucinate, we simplify the problem by considering two response values: whether $\mathit{q_{0}}$ produces any hallucination or not. Thus, we define the binary values for the probability of any hallucination as $p_{b}(q_{0})$.
$p_{b}\bigl{(}q_{0}\bigr{)}=\begin{cases}1&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}>0\\ 0&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}=0\\ \end{cases}$
Multi-class Labels
4 Metrics for Hallucination
Two key metrics are observed: accuracy and agreement. The concept of truthful hallucination focuses on the accuracy of individual agents. In addition, we analyze the level of agreement as an indicator of correlation. Moreover, the simulation’s reliability is assessed at the corpus level.
Accuracy
Accuracy serves as the measure of correctness in our evaluation by comparing the generated output $a_{i}$ to the ground truth string $y$, aiming for an almost exact match. For Multiple Choice queries, we also consider the choice label. However, due to minor variations in special characters, we adopt a partial, case-insensitive matching approach using the TheFuzz library. If there is a no match between the output $a_{i}$ and the ground truth $y$, we assign $\mathbb{I}\bigl{[}a_{i}\not=y\bigr{]}\mapsto 1$; otherwise, $\mathbb{I}\bigl{[}a_{i}=y\bigr{]}\mapsto 0$. To assess the range of correctness for each experiment, we compare the results of the original query $q_{0}$ and its output $a_{0}$ juxtaposed to the mode (most common) vote. In addition, we present the lower bound (all correct) and upper bound (at least one correct) accuracy across our $n+1=6$ agents.
Agreement
Accuracy alone is insufficient for evaluating the agreement among multiple agents when analyzing truthful hallucinations. To address this limitation, we utilize several statistical measures: Item Difficulty ($\mathbf{\mu_{D}}$) (Lord, 1952), Fleiss’s Generalized $\mathbf{\kappa}$ (Cohen, 1960; Fleiss, 1971), Mean Certainty ($\mathbf{H_{\eta}}$) (Shannon, 1948; Wilcox, 1973), and Gibbs’ $\mathbf{M_{2}}$ Index (Gibbs and Poston, 1975). These measures enable the assessment with regards to the level of agreement among independent samplings of a query, regardless of correctness. For example, if all LLM agents provide the same incorrect answer, high agreement indicates a misconception. Appendix D.3 expands in detail.
@@ 4 Metrics for Hallucination
Expected Rate Train Val Test
0.0% $(y=0)$ 139,123 17,146 9,202
16.7% $(y=1)$ 35,114 4,974 2,757
33.3% $(y=2)$ 20,213 3,371 1,967
50.0% $(y=3)$ 15,749 2,757 1,768
66.7% $(y=4)$ 14,477 2,735 1,970
83.3% $(y=5)$ 17,123 3,242 2,171
100.0% $(y=6)$ 60,693 10,266 3,019

### v3

# Is There No Such Thing As A Bad Question? H4R: HalluciBot For Ratiocination, Rewriting, Ranking, and Routing

@@ Abstract
Hallucination continues to be one of the most critical challenges in the institutional adoption journey of Large Language Models (LLMs). While prior studies have primarily focused on the post-generation analysis and refinement of outputs, this paper centers on the effectiveness of queries in eliciting accurate responses from LLMs. We present HalluciBot, a model that estimates the query’s propensity to hallucinate before generation, without invoking any LLMs during inference. HalluciBot can serve as a proxy reward model for query rewriting, offering a general framework to estimate query quality based on accuracy and consensus. In essence, HalluciBot investigates how poorly constructed queries can lead to erroneous outputs - moreover, by employing query rewriting guided by HalluciBot’s empirical estimates, we demonstrate that $95.7\%$ output accuracy can be achieved for Multiple Choice questions. The training procedure for HalluciBot consists of perturbing 369,837 queries $n$ times, employing $n+1$ independent LLM agents, sampling an output from each query, conducting a Multi-Agent Monte Carlo simulation on the sampled outputs, and training an encoder classifier. The idea of perturbation is the outcome of our ablation studies that measures the increase in output diversity ($+12.5$ agreement spread) by perturbing a query in lexically different but semantically similar ways. Therefore, HalluciBot paves the way to ratiocinate ($76.0\%$ test F1 score, $46.6\%$ in saved computation on hallucinatory queries), rewrite ($+30.2\%$ positive class transition from hallucinatory to non-hallucinatory), rank ($+50.6\%$ positive class transition from hallucinatory to non-hallucinatory), and route queries to effective pipelines.
Extended versionhttps://arxiv.org/abs/2404.12535
@@ 1 Introduction
Figure 1: Comparison of traditional inference methods and HalluciBot’s use-cases. In the former, the user inputs a query either through a direct inference or a retrieval-augmented generation (RAG) pipeline. If the output is hallucinatory, the user must decide whether to end the session or revise the query for successive generation rounds. In contrast, HalluciBot can be used to assess the query’s quality before generation. Therefore, users can gain insight into the hallucination risk (“Ratiocinate”), automate the query rewriting stage through informed feedback (“Rewrite”) or Best-of-N sampling across multiple candidates (“Rank”), and route the query across different operating modes (“Route”), since HalluciBot is scenario-aware (Extractive / Abstractive), potentially bypassing computationally expensive stages, such as RAG or Rewrite.
Despite the promising potential for a myriad of use cases, Large Language Models (LLMs) offer limited insights into their chain of thought (Liang et al. 2022; Wei et al. 2023; Kojima et al. 2023; Li et al. 2023) and have the propensity to hallucinate in various circumstances (Jiang et al. 2021). Common factors that drive hallucinations encompass high model complexity, flawed data sources, or inherent sampling randomness. Specifically, the intrinsic trade-off between greedy deterministic decoding and the creativity spawned through nucleus sampling induces a heightened propensity to hallucinate (Huang et al. 2023) - LLMs frequently advance output quality through different sampling methods (Holtzman et al. 2020; Fan, Lewis, and Dauphin 2018; Holtzman et al. 2018; Radford et al. 2018). The challenge of understanding hallucinations is compounded by limitations such as the frequent inaccessibility into the LLMs’ training datasets (Liang et al. 2022). HuggingFace’s release of its “Hallucinations Leaderboard” on January 29th, 2024 (Minervini et al. 2024; Gao et al. 2023) highlights the importance of resolving hallucination-related issues via the concerted effort of evaluating different LLMs. In this context, the majority of current studies have focused on the post-generation phase of output analysis as expanded in Peng et al. (2023) such as - (1) self-refinement via feedback loops on the model’s output (Madaan et al. 2023), (2) analysis of logit output values to detect hallucination (Varshney et al. 2023), or (3) for a minority of studies focused on the pre-generation phase, the ingestion of recent knowledge to improve performance (Tonmoy et al. 2024). We propose a novel model, HalluciBot, that predicts the probability of hallucination, before any generation, for a given query. In essence, this paper refocuses the study of hallucination to an empirical evaluation of the input query - how much does the query’s quality influence the model’s propensity to hallucinate? Therefore, HalluciBot estimates,
@@
    a binary classification of the query’s propensity to hallucinate (“Yes” or “No”), as well as,
@@
    a non reinforcement-learning method to guide query rewriting, enabling the construction of this encoder to be agnostic to closed-source or open-source LLMs (Ma et al. 2023).
We train HalluciBot as a binary classifier to predict whether a query will lead to erroneous outputs. To generate ground truth labels, we use a Multi-Agent Monte Carlo simulation that perturbs the query and checks for inaccuracies. If any perturbed version causes an error, the original query is labeled as hallucinatory. In this paper, HalluciBot leverages gpt-3.5-turbo, trained via (1) perturbing 369,837 queries $n$ times to retain the original semantic meaning yet diverge lexically, (2) employing $n+1$ independent agents to sample an output from each perturbation including the original query, at a temperature of $1.0$ for diversity, (3) conducting a Monte Carlo simulation on 2,219,022 sampled outputs, and (4) deriving an empirical estimate into the expected rate of hallucination $p_{h}(q_{0})$ for the original query. We prove that introducing perturbations before sampling $n+1$ outputs for query $q_{0}$ garners a $13.2$ point spread in the lower and upper bound accuracy, with a $12.5$ point decrease in Fleiss’s $\kappa$ for agreement, even as the modal accuracy remains largely unchanged ($1.3$ point difference). In other words, perturbations introduce more variability in the outputs, while preserving the central tendency. As HalluciBot generates the probability of hallucination in the pre-generation phase, the estimates can be used in a myriad of downstream modalities (Figure 1) such as: “Ratiocinate” to purely estimate the query’s quality; “Rewrite” to leverage the probabilities and improve the query’s quality via iterative feedback; “Rank” to rank perturbations, using probabilities as a proxy reward model in Best-of-N sampling; “Route” to route the best next steps, depending on scenarios such as Extractive or Abstractive. By cross-tabulating the predicted hallucination rates across scenarios, HalluciBot can act as a router, through which certain queries can be guided to a black-box LLM, while others will require a more complex pipeline including context retrieval, web search, or agents (Watson et al. 2023; Zeng et al. 2024; Cho et al. 2024).
Contributions. As a result, our study has culminated in the following pillars of contribution. (1) HalluciBot is the first encoder-based model to derive, before generation, an anticipated rate of hallucination for any type of query, achieving a validation accuracy of $73.6\%$ ($80.2\%$ F1) and a testing accuracy of $69.5\%$ ($76.0\%$ F1). (2) our approach to construct HalluciBot absorbs the computational complexity of Monte Carlo sampling, exploration, and training prior to the user session. Thus, institutions that employ HalluciBot can systematically save on the considerable amount of computational waste engendered by “highly probable” hallucinatory queries ($46.6\%$ in saved computation during testing). (3) the hallucination probability can be leveraged as a proxy reward model in a myriad of different infrastructures; HalluciBot paves the way to rewrite ($+31.9\%$ positive class transition from hallucinatory to non-hallucinatory), rank ($+51.4\%$ positive class transition from hallucinatory to non-hallucinatory), and route ($+60.0\%$ diverted) queries. (4) HalluciBot generalizes to systems with RAG or few-shot question answering systems with an LLM generator by differentiating the scenario in its prompt. Also, it can generalize to closed systems only accessible via API calls (OpenAI 2022; Google 2023; Microsoft 2023). (5) HalluciBot’s training methodology can be leveraged for any model or training corpus; it can be leveraged as a general means by which the research community can develop an encoder to assess query quality.
Figure 2: Training Overview. A single query, $q_{0}$, is perturbed in $n$ different ways. Next, The original and perturbed queries $q_{i}$ are independently answered by the Generator agents. This Multi-Agent Monte Carlo simulation provides an estimate into the rate of hallucination $p_{h}(q_{0})$ for the original query $q_{0}$. Via these simulated results, HalluciBot is trained to predict the probability that any $q_{0}$ could hallucinate, and predict the expected consensus of sampled outputs before generation.
@@ 2 Related Work
With regards to hallucination mitigation studies, an overwhelming majority focuses on the post-generation stage of analyzing outputs. A minority concentrates on the pre-generation phase and even amongst those, the focus lies in incorporating recent knowledge into LLMs. In detail, many expand on the universally utilized method of context-based retrieval systems (Reimers and Gurevych 2019; Johnson, Douze, and Jégou 2019; Nogueira and Cho 2020; Karpukhin et al. 2020; Lewis et al. 2020; Izacard and Grave 2021). Other methods include relying on the model’s general knowledge (Khashabi et al. 2020) or conditioning the QA model on context generated by the LLM itself (Yu et al. 2023). Certain work has focused on mitigating hallucinations by augmenting the way LLMs generate their answers. One of the more popular techniques is to have the model enumerate its chain-of-thought (Wei et al. 2023) and think step by step (Nye et al. 2021), while building context. Another method to augment generation with context is by semantic retrieval (Lewis et al. 2020; Liu et al. 2021), handling hallucinations as they arise (Varshney et al. 2023), or using LLMs to generate context rather than retrieve (Yu et al. 2023). PromptChainer (Wu et al. 2022) profiled techniques to craft LLM chains, in which the output of one LLM’s generation process, when fed into the next LLM, can allow for more complex tasks. Language Model Cascades (Dohan et al. 2022) demonstrated that LLMs can yield probabilistic programs to tackle multi-step reasoning problems. Self-consistency (Wang et al. 2023) leveraged a new decoding strategy to sample multiple generative pathways - then select the most consistent answer. Also, Kumar, Paria, and Tsvetkov (2022) explored gradient-based sampling procedures that satisfy user-defined constraints. Most recent work has focused on sampling-based calibration within a single model (Cole et al. 2023) or self-verification (Kadavath et al. 2022) - the latter focuses on generating a set of outputs and feeding those back into the LLM. Furthermore, Snyder, Moisescu, and Zafar (2023) explores how artifacts can differentiate hallucinated outputs. One common feature amongst these approaches is that the focus is on the output rather than the query. Alzahrani et al. (2024) explored how LLMs are highly sensitive to minute perturbations, such as changing the order of answer choices. Also, while Zheng and Saparov (2023) study lexical perturbations, no study on hallucinations employs a Multi-Agent approach coupled with query perturbations - which are hallmark features of HalluciBot.
@@
Extractive SQuADv2
Multiple Choice TruthfulQA, SciQ, MMLU, PIQA, BoolQ,
OpenBookQA, MathQA, ARC - E/C
Abstractive SQuADv2, TruthfulQA, SciQ,
@@
Figure 3: Distribution of the observed number of hallucinations per scenario. For Extractive, additional context mitigates the rate of hallucination. For Multiple Choice, distractors can cause confusion amongst agents uniformly. However, for Abstractive, no additional information can cause massive disparities in correctness - most of our simulations resulted in no or all hallucinations.
@@
Observed Rate Train Val Test
0.0% $(y=0/6)$ 139,123 17,146 9,202
16.7% $(y=1/6)$ 35,114 4,974 2,757
33.3% $(y=2/6)$ 20,213 3,371 1,967
50.0% $(y=3/6)$ 15,749 2,757 1,768
66.7% $(y=4/6)$ 14,477 2,735 1,970
83.3% $(y=5/6)$ 17,123 3,242 2,171
100.0% $(y=6/6)$ 60,693 10,266 3,019